# Adaptive-RAG: Learning to Adapt Retrieval-Augmented Large Language Models through Question Complexity

### 21 Mar 2024 v1

@@ 4 Experimental Setups
In this section, we explain datasets, models, metrics, and implementation details. We further provide the additional details in Appendix A.
@@ 4.1 Datasets
In order to simulate a realistic scenario, where different queries have varying complexities, in the unified experimental setting, we use both the single-hop and multi-hop QA datasets simultaneously.
Table 1: Averaged results on a collection of benchmark datasets for open-domain question answering including the single-hop and multi-hop queries, with different LLMs. Self-RAG is trained with the different base LLM, namely LLaMA2 Touvron et al. (2023); therefore, we compare the results of FLAN-T5-XL (3B) with the results from Self-RAG with LLaMA2 (7B) and the results of others with the results from Self-RAG with LLaMA2 (13B). We emphasize our results in bold, for easy comparisons.
@@ 4.2 Models
We compare our Adaptive-RAG against relevant models, including three retrieval-augmented LLM strategies (in Section 3.1) and the adaptive retrieval approaches Mallen et al. (2023); Asai et al. (2024), which can be grouped into one of three categories: Simple, Adaptive, and Complex. Specifically, Simple approaches include the 1) No Retrieval and 2) Single-step Approach-based methods. Adaptive approaches include the 3) Adaptive Retrieval Mallen et al. (2023), 4) Self-RAG Asai et al. (2024), and our 5) Adaptive-RAG, which can adaptively perform retrieval based on the question complexity. For the 6) Multi-step Approach, we use the most sophisticated state-of-the-art method Trivedi et al. (2023) iteratively accessing both the retriever and LLM with Chain-of-Thought reasoning Wei et al. (2022b), for every query. Note that models across different categories are not directly comparable. Yet, in the ideal setting, Adaptive approaches should be more effective than those in the Simple category while simultaneously being more efficient than the Complex one. Therefore, we also report the performance in an ideal scenario, 7) Adaptive-RAG w/ Oracle, using the oracle classifier with our Adaptive-RAG.
@@ 4.3 Evaluation Metrics
When it comes to evaluating adaptive models, it is essential to simultaneously consider both the task performance and efficiency along with their trade-offs. Thus, we report the results with five metrics, where three of them measure the effectiveness and the other two measure the efficiency. In particular, for effectiveness, we use F1, EM, and Accuracy (Acc) following the standard evaluation protocol Mallen et al. (2023); Baek et al. (2023); Asai et al. (2024), where F1 measures the number of overlapping words between the predicted answer and the ground truth, EM measures whether they are the same, and Acc measures whether the predicted answer contains the ground-truth answer. For efficiency, we measure the number of retrieval-and-generate steps and the average time for answering each query relative to the one-step approach.
@@ 5 Experimental Results and Analyses
First of all, Table 1 shows our main results averaged over all considered datasets, which corroborate our hypothesis that simple retrieval-augmented strategies are less effective than the complex strategy, while the complex one is significantly more expensive than the simple ones. In addition, we report the more granular results with FLAN-T5-XL on each of single-hop and multi-hop datasets in Table 2 (and more with different LLMs in Table 7 and Table 8 of Appendix), which are consistent with our hypothesis observed in Table 1.
However, in a real-world scenario, not all users ask queries with the same level of complexity, which emphasizes the importance of the need for adaptive strategies. Note that among the adaptive strategies, our Adaptive-RAG shows remarkable effectiveness over the competitors (Table 1). This indicates that merely focusing on the decision of whether to retrieve or not is suboptimal. Also, as shown in Table 2, such simple adaptive strategies are particularly inadequate for handling complex queries in multi-hop datasets, which require aggregated information and reasoning over multiple documents. Meanwhile, our approach can consider a more fine-grained query handling strategy by further incorporating an iterative module for complex queries. Furthermore, in a realistic setting, we should take into account not only effectiveness but also efficiency. As shown in Table 1, compared to the complex multi-step strategy, our proposed adaptive strategy is significantly more efficient across all model sizes. This is meaningful in this era of LLMs, where the cost of accessing them is a critical factor for practical applications and scalability. Finally, to see the upper bound of our Adaptive-RAG, we report its performances with the oracle classifier where the classification performance is perfect. As shown in Table 1 and Table 2, we observe that it achieves the best performance while being much more efficient compared to our Adaptive-RAG without the oracle classifier. These results support that our proposal for adapting retrieval-augmented LLM strategies based on query complexity is a valid and significant approach, and further suggests the direction to develop more improved classifies to achieve optimal performance.
@@
To understand how the proposed classifier works, we analyze its performance across different complexity labels. As Figure 3 (Left and Center) shows, the classification accuracy of our Adaptive-RAG is better than that of the other adaptive retrieval baselines, which leads to overall QA performance improvements. In other words, this result indicates that our Adaptive-RAG is capable of more accurately classifying the complexity levels with various granularities, which include not performing retrieval, performing retrieval only one time, and performing retrievals multiple times. In extension to the true positive performance of our classifier averaged over all those three labels in Figure 3 (Left and Center), we further report its confusion matrix in Figure 3 (Right). We note that the confusion matrix reveals some notable trends: ‘C (Multi)’ is sometimes misclassified as ‘B (One)’ (about 31%) and ‘B (One)’ as ‘C (Multi)’ (about 23%); ‘A (No)’ is misclassified oftentimes as ‘B (One)’ (about 47%) and less frequently as ‘C (Multi)’ (about 22%). While the overall results in Figure 3 indicate that our classifier effectively categorizes the three labels, further refining it based on misclassification results would be a meaningful direction for future work.
@@
While Table 1 shows the relative elapsed time for each of three different RAG strategies, we further provide the exact elapsed time per query for our Adaptive-RAG and the distribution for predicted labels from our query-complexity classifier in Table 3. Similar to the results of the elapsed time in Table 1 (relative time), Table 3 (exact time) shows that efficiency can be substantially improved by identifying simple or straightforward queries.
@@
We have shown that the classifier plays an important role in adaptive retrieval. Here, we further analyze the different strategies for training the classifier by ablating our full training strategy, which includes two approaches: generating silver data from predicted outcomes of models and utilizing inductive bias in datasets (see Section 3.2). As Table 4 shows, compared to the training strategy relying solely on the data derived from inductive bias, ours is significantly more efficient. This efficiency is partly because ours also takes into account the case that does not consider any documents at all, as also implied by the classification accuracy; meanwhile, queries in the existing datasets do not capture the information on whether the retrieval is not required. On the other hand, in the case of only using the silver data annotated from the correct predictions, while its overall classification accuracy is high, the overall QA performance implies that relying on the silver data may not be optimal. This may be because this silver data does not cover complexity labels on incorrectly predicted queries, which leads to lower generalization effect on queries relevant to them. Meanwhile, by also incorporating complexity labels from dataset bias (single-hop vs multi-hop), the classifier becomes more accurate in predicting multi-hop queries, leading to the better performance. It is worth noting that our automatic labeling strategies are two particular instantiations for training the classifier, and that there could be other instantiations, which we leave as future work.
@@
To investigate the sensitivity of our classifier according to its varying sizes, we conducted further experiments. As shown in Table 6, we observe no significant performance differences among classifiers of various sizes, even with reduced complexity and fewer parameters in smaller classifiers. This indicates that our proposed classifier can contribute to resource-efficient settings in real-use cases with smaller sizes without compensating performance.
@@
We conduct a case study to qualitatively compare our Adaptive-RAG against Adaptive Retrieval. Table 5 shows the classified complexity and the query handling patterns for both simple and complex questions. First, for the simple single-hop question, our Adaptive-RAG identifies that it is answerable by only the LLM’s parametric knowledge about ‘Google’. In contrast, Adaptive Retrieval fetches additional documents, leading to longer processing times and occasionally producing incorrect responses due to the inclusion of partially irrelevant information about ‘Microsoft’. Meanwhile, when faced with a complex question, our Adaptive-RAG seeks out relevant information, including details like ‘a son of John Cabot’, which may not have been stored in LLMs, while Adaptive Retrieval fails to request such information from external sources, resulting in inaccurate answers.
@@ 6 Conclusion
While our Adaptive-RAG shows clear advantages in effectiveness and efficiency by determining the query complexity and then leveraging the most suitable approach for tackling it, it is important to recognize that there still exists potential avenues for improving the classifier from the perspectives of its training datasets and architecture. Specifically, as there are no available datasets for training the query-complexity classifier, we automatically create new data based on the model prediction outcomes and the inductive dataset biases. However, our labeling process is one specific instantiation of labeling the query complexity, and it may have the potential to label queries incorrectly despite its effectiveness. Therefore, future work would create new datasets that are annotated with a diverse range of query complexities, in addition to the labels of question-answer pairs. Also, as the performance gap between the ideal classifier in Table 1 and the current classifier in Figure 3 indicates, there is still room to improve the effectiveness of the classifier. In other words, our classifier design based on the smaller LM is the initial, simplest instantiation for classifying the query complexity, and based upon it, future work may improve the classifier architecture and its performance, which will positively contribute to the QA performance.
@@
The experimental results on Adaptive-RAG validate its applicability in realistic scenarios, where a wide range of diverse user queries exists. Nonetheless, given the potential diversity of real-world user inputs, it is crucial to also consider scenarios where these inputs might be offensive or harmful. We should be aware that such inputs could lead to the retrieval of offensive documents and the generation of inappropriate responses by the retrieval-augmented LLMs. To address this challenge, developing methods to detect and manage offensive or inappropriate content in both user inputs and retrieved documents within the retrieval-augmented framework is essential. We believe that this is a critical area for future work.
@@ A.1 Datasets
We use publicly open datasets for both single-hop and multi-hop QA datasets, referring to Karpukhin et al. (2020) and Trivedi et al. (2023), respectively. We describe the characteristics of each dataset:
@@ A.2 Models
4) Self-RAG. This baseline Asai et al. (2024) trains the LLM to adaptively perform retrieval and generation, where the retrieval is conducted once it predicts the special retrieval token above a certain threshold, and the answer generation is followed.

### 28 Mar 2024 v2

@@ 4 Experimental Setups
In this section, we explain datasets, models, metrics, and implementation details. We provide additional details in Appendix A.
@@ 4.1 Datasets
In order to simulate a realistic scenario, where different queries have varying complexities, we use both the single-hop and multi-hop QA datasets simultaneously, in the unified experimental setting.
Table 1: Averaged results on a collection of benchmark datasets for open-domain question answering including the single-hop and multi-hop queries, with different LLMs. Self-RAG is trained with a different base LLM, namely LLaMA2 Touvron et al. (2023); therefore, we compare the results of FLAN-T5-XL (3B) with the results from Self-RAG with LLaMA2 (7B) and the results of others with the results from Self-RAG with LLaMA2 (13B). We emphasize our results in bold, for easy comparisons.
@@ 4.2 Models
We compare our Adaptive-RAG against relevant models, including three retrieval-augmented LLM strategies (in Section 3.1) and the adaptive retrieval approaches Mallen et al. (2023); Asai et al. (2024), which can be grouped into one of three categories: Simple, Adaptive, and Complex. Specifically, Simple approaches include the 1) No Retrieval and 2) Single-step Approach-based methods. Adaptive approaches include the 3) Adaptive Retrieval Mallen et al. (2023), 4) Self-RAG Asai et al. (2024), and our 5) Adaptive-RAG, which can adaptively perform retrieval based on the question complexity. For the 6) Multi-step Approach, we use the most sophisticated state-of-the-art method Trivedi et al. (2023), iteratively accessing both the retriever and LLM with Chain-of-Thought reasoning Wei et al. (2022b), for every query. Note that models across different categories are not directly comparable. Yet, in the ideal setting, Adaptive approaches should be more effective than those in the Simple category while simultaneously being more efficient than the Complex one. Therefore, we also report the performance in an ideal scenario, 7) Adaptive-RAG w/ Oracle, using the oracle classifier with our Adaptive-RAG.
@@ 4.3 Evaluation Metrics
When it comes to evaluating adaptive models, it is essential to simultaneously consider both the task performance and efficiency along with their trade-offs. Thus, we report the results with five metrics, where three of them measure the effectiveness and the other two measure the efficiency. In particular, for effectiveness, we use F1, EM, and Accuracy (Acc), following the standard evaluation protocol Mallen et al. (2023); Baek et al. (2023); Asai et al. (2024), where F1 measures the number of overlapping words between the predicted answer and the ground truth, EM measures whether they are the same, and Acc measures whether the predicted answer contains the ground-truth answer. For efficiency, we measure the number of retrieval-and-generate steps and the average time for answering each query relative to the one-step approach.
@@ 5 Experimental Results and Analyses
First of all, Table 1 shows our main results averaged over all considered datasets, which corroborate our hypothesis that simple retrieval-augmented strategies are less effective than the complex strategy, while the complex one is significantly more expensive than the simple ones. In addition, we report the more granular results with FLAN-T5-XL on each of the single-hop and multi-hop datasets in Table 2 (and more with different LLMs in Table 7 and Table 8 of Appendix), which are consistent with the results observed in Table 1.
However, in a real-world scenario, not all users ask queries with the same level of complexity, which emphasizes the importance of the need for adaptive strategies. Note that among the adaptive strategies, our Adaptive-RAG shows remarkable effectiveness over the competitors (Table 1). This indicates that merely focusing on the decision of whether to retrieve or not is suboptimal. Also, as shown in Table 2, such simple adaptive strategies are particularly inadequate for handling complex queries in multi-hop datasets, which require aggregated information and reasoning over multiple documents. Meanwhile, our approach can consider a more fine-grained query handling strategy by further incorporating an iterative module for complex queries. Furthermore, in a realistic setting, we should take into account not only effectiveness but also efficiency. As shown in Table 1, compared to the complex multi-step strategy, our proposed adaptive strategy is significantly more efficient across all model sizes. This is meaningful in this era of LLMs, where the cost of accessing them is a critical factor for practical applications and scalability. Finally, to see the upper bound of our Adaptive-RAG, we report its performances with the oracle classifier where the classification performance is perfect. As shown in Table 1 and Table 2, we observe that it achieves the best performance while being much more efficient than our Adaptive-RAG without the oracle classifier. These results support the validity and significance of our proposal for adapting retrieval-augmented LLM strategies based on query complexity, and further suggest the direction to develop more improved classifiers to achieve optimal performance.
@@
To understand how the proposed classifier works, we analyze its performance across different complexity labels. As Figure 3 (Left and Center) shows, the classification accuracy of our Adaptive-RAG is better than those of the other adaptive retrieval baselines, which leads to overall QA performance improvements. In other words, this result indicates that our Adaptive-RAG is capable of more accurately classifying the complexity levels with various granularities, which include not performing retrieval, performing retrieval only once, and performing retrieval multiple times. In addition to the true positive performance of our classifier averaged over all those three labels in Figure 3 (Left and Center), we further report its confusion matrix in Figure 3 (Right). We note that the confusion matrix reveals some notable trends: ‘C (Multi)’ is sometimes misclassified as ‘B (One)’ (about 31%) and ‘B (One)’ as ‘C (Multi)’ (about 23%); ‘A (No)’ is misclassified often as ‘B (One)’ (about 47%) and less frequently as ‘C (Multi)’ (about 22%). While the overall results in Figure 3 show that our classifier effectively categorizes the three labels, further refining it based on such misclassification would be a meaningful direction for future work.
@@
While Table 1 shows the relative elapsed time for each of the three different RAG strategies, we further provide the exact elapsed time per query for our Adaptive-RAG and the distribution for predicted labels from our query-complexity classifier in Table 3. Similar to the results of the elapsed time in Table 1 (relative time), Table 3 (exact time) shows that efficiency can be substantially improved by identifying simple or straightforward queries.
@@
We have shown that the classifier plays an important role in adaptive retrieval. Here, we further analyze the different strategies for training the classifier by ablating our full training strategy, which includes two approaches: generating silver data from predicted outcomes of models and utilizing inductive bias in datasets (see Section 3.2). As Table 4 shows, compared to the training strategy relying solely on the data derived from inductive bias, ours is significantly more efficient. This efficiency is partly because ours also takes into account the case that does not consider any documents at all, as also implied by the classification accuracy; meanwhile, queries in the existing datasets do not capture the information on whether the retrieval is required or not. On the other hand, in the case of only using the silver data annotated from the correct predictions, while its overall classification accuracy is high, the overall QA performance implies that relying on the silver data may not be optimal. This may be because this silver data does not cover complexity labels over incorrectly predicted queries, which leads to lower generalization effect on queries relevant to them. Meanwhile, by also incorporating complexity labels from dataset bias (single-hop vs multi-hop), the classifier becomes more accurate in predicting multi-hop queries, leading to the better performance. It is worth noting that our automatic labeling strategies are two particular instantiations for training the classifier, and that there could be other instantiations, which we leave as future work.
@@
To investigate the sensitivity of our classifier according to its varying sizes, we conducted further experiments. As shown in Table 6, we observe no significant performance differences among classifiers of various sizes, even with reduced complexity and fewer parameters in smaller classifiers. This indicates that our proposed classifier can contribute to resource-efficient settings in real-use cases with smaller sizes without compromising the performance.
@@
We conduct a case study to qualitatively compare our Adaptive-RAG against Adaptive Retrieval. Table 5 shows the classified complexity and the query handling patterns for both simple and complex questions. First, for the simple single-hop question, our Adaptive-RAG identifies that it is answerable by only using the LLM’s parametric knowledge about ‘Google’. By contrast, Adaptive Retrieval fetches additional documents, leading to longer processing times and occasionally producing incorrect responses due to the inclusion of partially irrelevant information about ‘Microsoft’. Meanwhile, faced with a complex question, Adaptive-RAG seeks out relevant information, including details like ‘a son of John Cabot’, which may not have been stored in LLMs, while Adaptive Retrieval fails to request such information from external sources, resulting in inaccurate answers.
@@ 6 Conclusion
While our Adaptive-RAG shows clear advantages in effectiveness and efficiency by determining the query complexity and then leveraging the most suitable approach for tackling it, it is important to recognize that there still exist potential avenues for improving the classifier from the perspectives of its training datasets and architecture. Specifically, as there are no available datasets for training the query-complexity classifier, we automatically create new data based on the model prediction outcomes and the inductive dataset biases. However, our labeling process is one specific instantiation of labeling the query complexity, and it may have the potential to label queries incorrectly despite its effectiveness. Therefore, future work may create new datasets that are annotated with a diverse range of query complexities, in addition to the labels of question-answer pairs. Also, as the performance gap between the ideal classifier in Table 1 and the current classifier in Figure 3 indicates, there is still room to improve the effectiveness of the classifier. In other words, our classifier design based on the smaller LM is the initial, simplest instantiation for classifying the query complexity, and based upon it, future work may improve the classifier architecture and its performance, which will positively contribute to the overall QA performance.
@@
The experimental results on Adaptive-RAG validate its applicability in realistic scenarios, where a wide range of diverse user queries exist. Nonetheless, given the potential diversity of real-world user inputs, it is crucial to also consider scenarios where these inputs might be offensive or harmful. We should be aware that such inputs could lead to the retrieval of offensive documents and the generation of inappropriate responses by the retrieval-augmented LLMs. To address this challenge, developing methods to detect and manage offensive or inappropriate content in both user inputs and retrieved documents within the retrieval-augmented framework is essential. We believe that this is a critical area for future work.
Acknowledgements
This work was supported by Institute for Information and communications Technology Promotion (IITP) grant funded by the Korea government (No. 2018-0-00582, Prediction and augmentation of the credibility distribution via linguistic analysis and automated evidence document collection), Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2023-00275747), and the Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City.
@@ A.1 Datasets
We use publicly open datasets for both single-hop and multi-hop QA datasets, referring to as Karpukhin et al. (2020) and Trivedi et al. (2023), respectively. We describe the characteristics of each dataset:
@@ A.2 Models
4) Self-RAG. This baseline Asai et al. (2024) trains the LLM to adaptively perform retrieval and generation, where the retrieval is conducted once it predicts the special retrieval token above a certain threshold, and the answer generation follows.